1 a) Find the area of the surface obtained by rotating the circle x^2 + y^2 = 49 about the line y=7. (Keep two decimal places) (note: the answer is not 6,770.55)

b) According to the National Health Survey, the heights of adult males in the United States are
(normally distributed with mean) 73 inches, and standard deviation of 2.8 inches. What is the
probability that an adult male chosen at random is between 71 inches and 75 inches tall? (note: the answer is not 44.7)

c) The standard deviation for a random variable with probability density function f and mean μ is
defined
0= [(2 – u)? f (a) der]
Find the standard deviation for an exponential density function with mean 10. (note: the answer is not 0.44)
Answer #1

Solutions Ta Given circle x?+y= 49 Cine y = 7 ie, about x-axis Surface area is given as s=Satyds where as = + (dy je da y = Ja) Pl71c4c75) = P((71–73)/2.8) <(x-)/6< (75-73)/2-8) = P(-0-7142.22 0.1142) = P (220-7142) - P(22-0-7142) 70.1142 e tzte 31 d© The standard deviation for a random probability density function fand mean Given mean = 10. variable with is defined as f (75 °*° (PA)au – su gºrf(x)dk + 1 2 1 2 Now Jº xf(x)dx = d. Jo fldx = 1 Sum of all probabilities = 1 Substituting in on = 10v2